A motor boat departs from the north bank of a river and bears south to the opposite bank. The speed of the boat in calm waters is 20 mi/h, and the river runs east at 3 mi/h. What is the actual speed of the boat and what is its final bearing? (20.2 mph, S 8.5° W)
Answer by ikleyn(52807)

.

To find the actual speed, apply the Pythagorean theorem.


To find the angle, use the tangent function.


--------------

Comment from student: asb=sqrt(20^2+3^2)=sqrt(400+9)=sqrt(409) ≈20.22 mph aSW=atan(3/20)=8.53°



My response : Perfect (!)

I am very impressed and glad to see how perfectly did you manage to solve this problem (!).

My congrats (!)